At the end of the 2000 WNBA regular season, the Houston Comets had 22 more victories than loses. The number of victories they had was three less than six times the number of losses. How many regular season games did the Houston Comets play during the 2000 WNBA season?
Answer by stanbon(75887)

At the end of the 2000 WNBA regular season, the Houston Comets had 22 more victories than loses. The number of victories they had was three less than six times the number of losses. How many regular season games did the Houston Comets play during the 2000 WNBA season?
--------
Equations:
v = l + 22
v = 6*l -3
-----
Substitute and solve for "l":
l + 22 = 6*l -3
5*L = 25
L = 5 (# of losses)
-----
v = L + 22
v = 27 (# of victories)
===========================
Cheers,
Stan H.
==============